OpenStudy (anonymous):

Find the Laplace Transform of the given function. Assume that term-by-term integration of the infinite series is permissible.

OpenStudy (anonymous):

\[\huge f(t)=1+\sum_{k=1}^{\infty}(-1)^ku_k(t)\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Should I begin by expanding the summation and then just find the Laplace Transform of the first few terms?

OpenStudy (astrophysics):

.

OpenStudy (anonymous):

Hmm. I think I can rewrite it as \[\mathcal{L}[f(t)]=\frac{1}{s}+\sum_{k=1}^{\infty}(-1)^k~\mathcal{L}[u_k(t)]\]But what to do with that, no clue..

OpenStudy (anonymous):

\[\mathcal{L}[f(t)]=\frac{1}{s}+\sum_{k=1}^{\infty}(-1)^k~\frac{e^{-cs}}{s}\]

OpenStudy (anonymous):

Maybe?

ganeshie8 (ganeshie8):

\[\mathcal{L}(f(t-a)u(t-a)) = e^{-as}F(s)\]

ganeshie8 (ganeshie8):

in your case \(f(t-a) = 1\)

ganeshie8 (ganeshie8):

must be a typo, you mean \[\mathcal{L}[f(t)]=\frac{1}{s}+\sum_{k=1}^{\infty}(-1)^k~\frac{e^{-\color{red}{k}s}}{s}\]

OpenStudy (anonymous):

Ah yes, yes X) oops..

ganeshie8 (ganeshie8):

looks good to me, lets check with wolfram

OpenStudy (anonymous):

How do I rid the sum?

OpenStudy (anonymous):

Hmm, but during the test I cannot check Wolfram =/ I managed to find the same question elsewhere, and they have the same steps I do until about halfway and then they start doing something I don't understand.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

hey

ganeshie8 (ganeshie8):

wolfram is just for double checking your answer so that you feel confident

ganeshie8 (ganeshie8):

it is just a geometric series : \[\sum\limits_{k=0}^{\infty} \clubsuit^k = \dfrac{1}{1-\clubsuit}\] for \(|\clubsuit|\lt 1\)

OpenStudy (anonymous):

ohh wow haha, thanks!